OpenStudy (anonymous):

Identify the axis of symmetry for the graph of g(x) = 12(x + 6)2 - 5

hero (hero):

Once again, the x axis of symmetry will be the x that makes the binomial (x+6)^2 = 0 the only x that works is x = -6

OpenStudy (anonymous):

thanks, hero.

hero (hero):

You can also do it the long way and transform the expression to 12x^2 + 144x + 427 and use x = -b/2a Either way, you still end up with x = -6